OpenStudy (tester97):

x^2+4x=5 how do i solve?

OpenStudy (charlotte123):

I use wolframalpha OwO http://www.wolframalpha.com/input/?i=x%5E2%2B4x%3D5 x = -5 and x = 1

OpenStudy (tester97):

no char how do i solve it myself?

OpenStudy (anonymous):

try guessing small integers for x and see what you get.

OpenStudy (tester97):

ehh nvm i got it

OpenStudy (shamil98):

I'll give you a detailed explanation. x^2+4x=5 you subtract from both sides the 5. x^2 + 4x - 5 = 0 Then you use the magical quadratic formula. \[x=\frac{ -b \pm \sqrt {b^2 -4(a)(c)} }{ 2(a) }\]

OpenStudy (shamil98):

Inputting your values of a, b, c. a = 1 , b = 4 , c = -5. \[x = \frac{ -4 \pm \sqrt { 16 + 20} }{ 2 } = \frac{ -4 \pm 6 }{ 2} \] \[\frac{ -4 + 6 }{ 2 } = 1\] \[\frac{ -4 - 6 }{ 2 } = -5\] your solutions are x = 1 , and x = -5

OpenStudy (charlotte123):

what sham said O_O